OpenStudy (rowanwhitethorn):

Need help with absolute value equation: 5| 2x-4| +1 =11

OpenStudy (otherworldly):

2x = 9

jhonyy9 (jhonyy9):

like a first step subtract from both sides 1 so what will get ?

jhonyy9 (jhonyy9):

please collaborate

OpenStudy (otherworldly):

Move all coefficients to the other side of the equal sign where the 11 is and change all the coeffecients signs when moving it to the opposite sign

jhonyy9 (jhonyy9):

5I2x+4I +1 =11 5I2x+4I +1 -1 = 11 -1 ?

jhonyy9 (jhonyy9):

continue it please

OpenStudy (rowanwhitethorn):

5| 2x+4| =10?

OpenStudy (otherworldly):

yes

jhonyy9 (jhonyy9):

yes sure right this is easy so now divide both sides by 5 what will get ?

OpenStudy (rowanwhitethorn):

x I 2x+4 l =2

jhonyy9 (jhonyy9):

why x in front of this line of absolute value ?

jhonyy9 (jhonyy9):

there was 5 what you simplified it dividing both sides by 5 - or not ?

OpenStudy (rowanwhitethorn):

opps sorry i meant 1l 2x+4l =2

jhonyy9 (jhonyy9):

without this 1 I2x+4I = 2